1) dx/dy+4y=12 for y(0)=1 (general nad particular solution)

1
Expert's answer
2021-12-20T16:43:34-0500

"\displaystyle
\frac{dx}{dy}=12-4y\\
dx= dy(12-4y)\\
x = 12y - 2y^2\\
\implies 12y -2y^2-x =0\\
y(0) =1\\
\implies 12-2-x =0\\
x=10\\
\text{The general solution is given by}\\
12y -2y^2-10=0"